OpenStudy (anonymous):

For a real number $\alpha > 0 $ call $\displaystyle \binom{\alpha}{n}:=\frac{\alpha(\alpha-1)(\alpha-2)...(\alpha+1-n)}{n!}$ the generalized binomial coeeficient. The products in the denominator and the numerator have $n$ factors for each. Assume that $\alpha \notin \mathbb{N}$.\\\\ a) Prove: $\lim_{n \to \infty}\frac{\binom{\alpha}{n+1}}{\binom{\alpha}{n}} = 1$

OpenStudy (paxpolaris):

\[\displaystyle \binom{\alpha}{n}:=\frac{\alpha(\alpha-1)(\alpha-2)...(\alpha+1-n)}{n!}\] Prove: \[\Large \lim_{n \to \infty}\frac{\binom{\alpha}{n+1}}{\binom{\alpha}{n}} = 1\]

OpenStudy (paxpolaris):

\[\Large {{\alpha \choose n+1} \over {\alpha \choose n}}={a-n \over n+1}\]

OpenStudy (anonymous):

ls It answer?

OpenStudy (paxpolaris):

I just simplified it.... you still need to take the limit: \[\Large \lim_{n \to \infty}{\alpha-n \over n+1}\]

OpenStudy (paxpolaris):

i don't know how it is =1 :'(

OpenStudy (anonymous):

thanks your effort tomorrow i must give very important homework i think i need a longer solution so that my mentor give me enough points :( and i have no idea how to start..

OpenStudy (anonymous):

\[(α(α−1)(α−2)...(α−n+1)(α−n)/(n+1)!)/(α(α−1)(α−2)...(α−n+1)/n!)\] \[=(α−n)/(n+1)!)/1/n!=(α−n)n!/(n+1)!=(α−n)/(n+1)=α/(n+1)-n/(n+1)\] When n goes to infinity, α/(n+1) goes to 0, and -n/(n+1) goes to -1. So I get -1. Let me try to do it again....

OpenStudy (anonymous):

ok it sounds hopefully

OpenStudy (kinggeorge):

I'm honestly not seeing how it could be something other than -1. If you take some number \(\alpha\) and fix it in \(\mathbb{R}^+ \setminus \mathbb{N}\), Then as \(n\) goes to infinity, there comes a point where\[\binom{\alpha}{n}>0\]\[\binom{\alpha}{n+1}<0\]Because of the \(\alpha\cdot(\alpha-1)\cdot...\cdot(\alpha-(n-1))\) terms.This means that the numerator and denominator of \[\Large {{\alpha \choose n+1} \over {\alpha \choose n}}\]will have differing signs when \(\alpha-n<0\) and \(\alpha-n+1>0\). Hence, as \(n\) goes to infinity, it must be negative. And from what nightwill just showed above, it should be exactly \(-1\).

OpenStudy (kinggeorge):

I was a little unclear at the end. They will have differing signs because if \(\alpha\cdot(\alpha-1)\cdot...\cdot(\alpha-(n+1))\) has an odd number of negative numbers, then \(\alpha\cdot(\alpha-1)\cdot...\cdot(\alpha-n)\) must have an even number of negative numbers and vice versa as \(\displaystyle n\rightarrow \infty\)

OpenStudy (anonymous):

Also KongGeorge are you agreed with the solution of Nightwill?

OpenStudy (kinggeorge):

I'm agreeing with nightwill

OpenStudy (anonymous):

ok thank you very much for your efforts

OpenStudy (kinggeorge):

Also, @PaxPolaris take a look at http://www.wolframalpha.com/input/?i=6.5+choose+9 The property you listed only seems to hold when \(\alpha\) is a natural number.

OpenStudy (paxpolaris):

right sorry...